# Sustainable Greenhouse Microclimate Modeling: A Comparative Analysis of Recurrent and Graph Neural Networks

### v3

@@ Abstract
The integration of photovoltaic (PV) systems into greenhouses not only optimizes land use but also enhances sustainable agricultural practices by enabling dual benefits of food production and renewable energy generation. However, accurate prediction of internal environmental conditions is crucial to ensure optimal crop growth while maximizing energy production. This study introduces a novel application of Spatio-Temporal Graph Neural Networks (STGNNs) to greenhouse microclimate modeling, comparing their performance with traditional Recurrent Neural Networks (RNNs). While RNNs excel at temporal pattern recognition, they cannot explicitly model the directional relationships between environmental variables. Our STGNN approach addresses this limitation by representing these relationships as directed graphs, enabling the model to capture both environmental dependencies and their directionality. Using high-frequency data collected at 15-minute intervals from a greenhouse in Volos, Greece, we demonstrate that RNNs achieve exceptional accuracy in winter conditions ($R^{2}=0.985$) but show limitations during summer cooling system operation. Though STGNNs currently show lower performance (winter $R^{2}=0.947$), their architecture offers greater potential for integrating additional variables such as PV generation and crop growth indicators.
@@ 1 Introduction
Modern agriculture faces increasing pressure to optimize land use while reducing its environmental impact. The integration of photovoltaic (PV) systems into agricultural greenhouses has emerged as a promising solution to these challenges, offering dual benefits of food production and renewable energy generation. However, this integration introduces new complexities in greenhouse management, as PV panels directly influence light distribution, temperature patterns, and overall microclimate dynamics. These Agri-PV systems optimize land use by combining solar energy harvesting with protected crop cultivation, offering potential benefits for both energy and food security. Yet the successful implementation of PV-integrated greenhouses requires accurate prediction and control of the internal microclimate to ensure optimal growing conditions while maximizing energy production. While sophisticated physical models incorporating greenhouse thermodynamics, crop physiology, and PV system performance exist, their complexity often restricts real-time applications and rapid design optimization. This study, conducted within the framework of the European REGACE project, represents an initial step toward developing a comprehensive digital twin for PV-integrated greenhouses. Rather than relying on complex physical models, we explore empirical approaches that employ advanced machine learning to predict greenhouse microclimates. The focus on neural network-based models aims to lay the foundation for a more complete digital twin that will ultimately integrate crop growth dynamics and PV energy production. By comparing traditional Recurrent Neural Networks (RNNs) with directed Spatio-Temporal Graph Neural Networks (STGNNs), we investigate whether the explicit modeling of environmental relationships and directional influences can improve prediction accuracy while maintaining computational efficiency. Effective control and optimization of PV-integrated greenhouse environments present a fundamental challenge. Accurate climate control software integrating real-time data from environmental control computers (ECCs) can significantly enhance greenhouse management (Aaslyng et al., 2005). Nonetheless, internal parameters such as temperature and humidity are influenced by various external conditions, including ambient temperature, solar radiation, and wind speed. Conventional modeling approaches can fail to capture the nonlinear and dynamic nature of these relationships, especially when greenhouse data are collected at high frequency (Aaslyng et al., 2003). Researchers have therefore pursued increasingly sophisticated methods to balance these complexities, from physical simulations to data-driven models. Greenhouses represent intricate thermodynamic systems where external conditions and the internal microclimate interplay to influence crop growth and energy use. The introduction of PV panels adds a layer of complexity, prompting further advancements in modeling techniques (Ouazzani Chahidi et al., 2021; Baglivo et al., 2020; Brækken et al., 2023). These approaches range from detailed physics-based studies focusing on ventilation and evapotranspiration (Stanciu et al., 2016; Abdel-Ghany & Kozai, 2006; Mobtaker et al., 2019; Singh et al., 2018; Fitz-Rodríguez et al., 2010) to data-driven or hybrid solutions capable of simulating both microclimate and crop growth processes (Massa et al., 2011). While such models provide essential insights into greenhouse behavior, they can be computationally intensive, and their parameterization often demands multidisciplinary expertise. To address these constraints, researchers have increasingly explored empirical, data-driven approaches that learn greenhouse behavior directly from observations. Early neural network applications demonstrated promise (Seginer et al., 1994; Bussab et al., 2007), evolving into more sophisticated architectures such as Radial Basis Function Neural Networks (Ferreira et al., 2002) and optimized Multi-layer Perceptrons (D’Emilio et al., 2012). A major leap forward came with Recurrent Neural Networks (RNNs), notably Long Short-Term Memory (LSTM) variants, which excel at modeling time-dependent relationships in greenhouse environments (Fourati & Chtourou, 2007; Gharghory, 2020; Hongkang et al., 2018). In parallel, Nonlinear AutoRegressive models with eXogenous inputs (NARX) have shown effectiveness in capturing complex greenhouse system dynamics (Manonmani et al., 2018; Gao et al., 2023). Despite these advances, conventional neural networks predominantly treat environmental variables independently, limiting their ability to depict spatial or directional interconnections. This gap becomes especially relevant in PV-integrated greenhouses, where factors such as solar radiation, humidity, and temperature form intricate environmental interactions that purely temporal models may overlook. Spatio-Temporal Graph Neural Networks (STGNNs) address this shortfall by explicitly modeling intervariable relationships through graph structures (Yu et al., 2018; Sun et al., 2022). Here, each environmental variable becomes a node in a directed graph, capturing how changes in one parameter can propagate to others. The Graph Attention Network (GAT) framework further refines this approach by dynamically weighting node and edge influences, enabling simultaneous modeling of environmental dependencies and temporal evolution (Veličković et al., 2018). Such techniques align with the concept of digital twins in greenhouse systems, where software-based integrative models can guide real-time climate strategies (Aaslyng et al., 2005, 2003) and nutrient management (Massa et al., 2011).
In this study we apply and compare RNNs and directed STGNNs for greenhouse temperature prediction under multiple seasonal conditions. Using data from a Mediterranean greenhouse in Volos, Greece, we assess each model’s predictive accuracy, interpretability via attention mechanisms, and computational feasibility. By evaluating whether directional modeling improves performance over simpler temporal approaches, our work advances the broader REGACE objective of developing efficient, sustainable agrivoltaic greenhouses that integrate microclimate modeling with PV energy production and crop modeling.
@@ 2 Data
The data used in this study originates from a PV greenhouse located in Volos, Greece, for the year 2020, covering from January 9, 2020, to December 21, 2020. The dataset consists of measurements recorded at 15-minute intervals, resulting in 33,446 data points.
@@
The graphs of the temperature inside and outside the greenhouse for the three periods considered follow:
Figure 1: Temperatures inside and outside the greenhouse for the summer period
Figure 2: Temperatures inside and outside the greenhouse for the autumn period
Figure 3: Temperatures inside and outside the greenhouse for the winter period
To capture the complex relationships between these environmental factors, we represent them as nodes in a directed graph, where edges indicate the direction of influence from one variable to another. The graph is presented in Figure 4.
@@
These physical relationships form the basis for the graph structure shown in Figure 4, where each edge represents a direct causal influence between variables. This directed graph approach allows our STGNN model to learn and weight these physical dependencies during the prediction process.
Figure 4: Features’ interaction graph
Moreover, the dataset was divided into three seasons to better investigate the seasonal variations in the greenhouse microclimate and due to the high number of missing data. The dates of the divisions are:
@@
    Summer111 During the summer period, the greenhouse employs a cooling system to prevent internal temperatures from rising excessively.: From July 1 to August 31,
@@
    Autumn: From October 9 to November 30,
  • 3.
    Winter: From January 9 to February 29.
We used seasonal names for these periods, even though the dates do not perfectly correspond to the year seasons. This discrepancy is due to a significant number of missing values. For the same reason, spring data were excluded from the analysis due to many missing values recorded between April 23 and April 27.
2.1 Data handling
2.1.1 Missing values
Table 1 shows the number of missing values for each variable in each season.
Table 1: Number of missing values for each season
Season G2_temp G2_RH OUT_temp OUT_RH OUT_rad wind_speed
Summer 0 0 0 0 0 0
Winter 68 68 2 2 2 2
Autumn 75 75 75 75 75 75
To address missing data, we employed a combination of linear interpolation and rolling mean techniques. Linear interpolation is effective for high-density time series data, as it estimates missing values by creating a linear equation between the nearest known data points (Malvar et al., 2004; Taewon et al., 2019), as in Equation 1:
$f(x)=f(x_{o})+\frac{f(x_{1})-f(x_{o})}{x_{1}-x_{o}}(x-x_{o})$ (1)
where $f$ is the linear interpolation function, $x$ the target value and the subscripts 00 and $1$ represent respectively the previous and next value from target value. This approach ensured that the imputed values preserved the continuity and trends in the time series without introducing noticeable distortions or flattening in the data sequences. The rolling mean technique was applied in cases where data exhibited more noise or short-term fluctuations, allowing us to smooth these variations while still accounting for missing values.
2.1.2 Normalization
All features were normalized using the min-max method. This scaling brings all values into the range $[0,1]$, bringing all the variables expressed in different units to the same scale and making them comparable (Alaimo & Seri, 2021).
2.1.3 Train-Test Split
To prepare the data for modeling, each seasonal subset of the dataset was divided into training and testing sets while preserving the chronological order. This approach ensured that the integrity of the time series analysis was maintained, as no future information was inadvertently introduced into the training process. The first 80% of observations within each season were allocated to the training set, while the remaining 20% were used for testing. The resulting date ranges for the training and testing sets in each seasonal period are as follows:
  • 1.
    Winter:
    • (a)
      Training: from 2020-01-09 16:30:00 to 2020-02-19 04:00:00
    • (b)
      Testing: from 2020-02-19 04:15:00 to 2020-02-29 07:00:00
  • 2.
    Summer:
    • (a)
      Training: from 2020-07-01 14:30:00 to 2020-08-19 02:45:00
    • (b)
      Testing: from 2020-08-19 03:00:00 to 2020-08-31 06:00:00
  • 3.
    Autumn:
    • (a)
      Training: from 2020-10-09 22:00:00 to 2020-11-20 13:00:00
    • (b)
      Testing: from 2020-11-20 13:15:00 to 2020-11-30 22:45:00
2.2 Stationarity
In order to make statistical inference based on time series data, it is essential to ensure that the underlying data generative process fulfill the stationary property (Kirchgässner et al., 2012). Stationarity means that the statistical properties of the time series, such as mean and variance, remain consistent over time. This consistency is crucial because it ensures that the patterns observed in the past data are applicable to the future. A time series is stationary if it meets two conditions:
  1. 1.
    The mean (average value) of the series is constant over time.
  2. 2.
    The covariance between two points depends only on the time interval between them, not on their specific positions in time.
Stationarity implies that the series fluctuates around a constant level with consistent variation. This property is essential for making reliable inferences and predictions based on the data (Tsay, 2005). To verify that our data met these criteria, we applied the augmented Dickey-Fuller (ADF) test (Cheung & Lai, 1995). The ADF test results indicate that for all variables in all the three considered periods, the null hypothesis of non-stationarity is rejected, confirming that the series are stationary. The results of the ADF test are presented in Tables 4, 5 and 6 in Appendix A.2.1.
@@ 3 Methodology
Neural Networks (NNs) are computational models inspired by the human brain’s neural structure. They consist of interconnected nodes (neurons) organized into layers: an input layer, one or more hidden layers, and an output layer. Each connection between neurons has an associated weight that is adjusted during training to minimize prediction error. Deep Neural Networks (DNNs) extend NNs by incorporating multiple hidden layers, enabling the modeling of complex, nonlinear relationships in data. Activation functions such as ReLU, sigmoid, and tanh introduce nonlinearity, while backpropagation adjusts the weights based on the error rate obtained in previous epochs (Nielsen, 2015).
@@ 3.1 Recurrent neural networks
Recurrent Neural Networks (RNNs) are specialized neural networks designed to process sequential data, making them highly suitable for time-series analysis (Jain & Medsker, 1999). Unlike traditional feedforward neural networks, RNNs have recurrent connections that allow information to persist across time steps, enabling the network to capture temporal dependencies.
@@
$\hat{y}_{t}=\bm{W}_{y}\bm{h}_{t}+\bm{b}_{y}$
where $\bm{W}_{h}$, $\bm{U}_{h}$, and $\bm{W}_{y}$ are weight matrices. $\bm{b}_{h}$ and $\bm{b}_{y}$ are bias vectors. $\tanh$ is the hyperbolic tangent activation function introducing nonlinearity. In deeper RNNs, multiple RNN layers are stacked, passing the hidden state from one layer to the next, allowing the model to capture more complex temporal patterns (Nketiah et al., 2023).
RNNs are effective for modeling how current and past environmental conditions influence future states, making them ideal for predicting variables like greenhouse temperature (Rodriguez et al., 1999). A known challenge with RNNs is the vanishing gradient problem, where gradients diminish during backpropagation through time, making it difficult to learn long-term dependencies (Behrang et al., 2010). However, since our data exhibits short-term dependencies, this issue do not impact our study (Dubinin & Effenberger, 2024).
@@ 3.1.1 RNN implementation
We begin by preparing the dataset, where each input sequence $\bm{X}^{(i)}$ consists of historical environmental measurements over a period of $T$ time steps. The features include greenhouse temperature and relative humidity, as well as outdoor environmental variables such as temperature, relative humidity, solar radiation, and wind speed. The target value $y^{(i)}$ is the greenhouse temperature at the next time step.
@@
The model is compiled using the Adam optimizer (Kingma & Ba, 2014), which adapts the learning rate during training for efficient convergence. The loss function is set to Mean Squared Error (MSE), suitable for regression tasks.
@@ 3.2 Spatio-Temporal Graph Neural Networks
To model the spatio-temporal dynamics of the greenhouse microclimate, we employ directed STGNN. This approach integrates Graph Attention Networks (GATs) for spatial feature extraction (Veličković et al., 2018) and Long Short-Term Memory (LSTM) networks (Hochreiter & Schmidhuber, 1997) for temporal sequence modeling. Unlike traditional RNN-based models, which focus primarily on temporal sequences, STGNNs leverage Graph Neural Networks (GNNs) to incorporate topological and relational information among variables. This is well-aligned with recent advances in GNNs for time series (Jin et al., 2024). By representing each environmental variable, such as internal temperature, humidity, and external factors like solar radiation and wind speed, as a node in a directed graph, we capture how changes in one variable may propagate to others through directional edges. Figure 4 illustrates this architecture. Each node corresponds to an environmental variable, while edges encode the direction of influence. At each time step $t$, node features $\bm{X}_{t}$ are extracted from the input sequence $\bm{X}^{(i)}$. To prevent data leakage, the feature of the target node (greenhouse temperature) at the current time step is masked (set to zero) before model input. The GAT layers learn to weigh the importance of edges and neighboring nodes, revealing spatial dependencies, while the LSTM layers capture long-term temporal patterns.

### v4

@@ Abstract
The integration of photovoltaic (PV) systems into greenhouses not only optimizes land use but also enhances sustainable agricultural practices by enabling dual benefits of food production and renewable energy generation. However, accurate prediction of internal environmental conditions is crucial to ensure optimal crop growth while maximizing energy production. This study introduces a novel application of Spatio-Temporal Graph Neural Networks (STGNNs) to greenhouse microclimate modeling, comparing their performance with traditional Recurrent Neural Networks (RNNs). While RNNs excel at temporal pattern recognition, they cannot explicitly model the directional relationships between environmental variables. Our STGNN approach addresses this limitation by representing these relationships as directed graphs, enabling the model to capture both environmental dependencies and their directionality. We benchmark RNNs against directed STGNNs on two 15-min-resolution datasets from Volos (Greece): a six-variable 2020 installation and a more complex eight-variable greenhouse monitored in autumn 2024. In the simpler 2020 case the RNN attains near-perfect accuracy, outperforming the STGNN. When additional drivers are available in 2024, the STGNN overtakes the RNN ($R^{2}=0.905$ vs $0.740$), demonstrating that explicitly modelling directional dependencies becomes critical as interaction complexity grows. These findings indicate when graph-based models are warranted and provide a stepping-stone toward digital twins that jointly optimise crop yield and PV power in agrivoltaic greenhouses.
@@
{highlights}
Compare RNNs and directed STGNNs for greenhouse microclimate prediction.
RNN excels in the simple 2020 scenario.
STGNN beats RNN on a complex 2024 greenhouse with eight environmental drivers.
Directional graphs help when added variables amplify interaction paths.
Result guides when to pick temporal vs. spatio-temporal models.
@@ 1 Introduction
Modern agriculture faces increasing pressure to optimize land use while reducing its environmental impact. The integration of photovoltaic (PV) systems into agricultural greenhouses has emerged as a promising solution to these challenges, offering dual benefits of food production and renewable energy generation. However, this integration introduces new complexities in greenhouse management, as PV panels directly influence light distribution, temperature patterns, and overall microclimate dynamics. These Agri-PV systems optimize land use by combining solar energy harvesting with protected crop cultivation, offering potential benefits for both energy and food security. Yet the successful implementation of PV-integrated greenhouses requires accurate prediction and control of the internal microclimate to ensure optimal growing conditions while maximizing energy production. While sophisticated physical models incorporating greenhouse thermodynamics, crop physiology, and PV system performance exist [1], their complexity often restricts real-time applications and rapid design optimization. This study, conducted within the framework of the European REGACE project, represents an initial step toward developing a comprehensive digital twin for PV-integrated greenhouses. Rather than relying on complex physical models, we explore empirical approaches that employ advanced machine learning to predict greenhouse microclimates. By comparing traditional Recurrent Neural Networks (RNNs) with directed Spatio-Temporal Graph Neural Networks (STGNNs), we investigate whether explicit modeling of environmental relationships and directional influences can improve prediction accuracy while maintaining computational efficiency.
RNNs capture temporal patterns but treat each variable independently, whereas STGNNs embed each environmental variable as a node in a directed graph, allowing the model to learn how changes propagate through explicit causal pathways [2, 3, 4]. We analyse 15-min data from two Mediterranean greenhouses in Volos (Greece): GH2 (2020, six variables, $\approx$33 k samples across three seasons) and the more complex GH4 (autumn 2024, eight variables including PAR and CO2, 4 k training / 1 k test). Both models use identical hyperparameter budgets and a 96-step (24 h) input window, ensuring that performance differences arise from the modelling paradigm rather than tuning effort. A vanilla RNN achieves near-perfect accuracy on GH2, yet the STGNN overtakes it on GH4 ($R^{2}=0.905$ vs 0.740), demonstrating that directional graphs earn their keep when additional sensors amplify interaction pathways. This finding addresses the open question of when graph machinery is necessary for greenhouse prediction, complementing earlier empirical studies that relied solely on temporal models [5, 6, 7, 8, 9, 10, 11, 12]. It also advances digital-twin research for PV-integrated houses by showing how physics-informed edge sets can improve data-driven predictions without exhaustive search. Effective micro-climate control remains challenging because internal temperature and humidity respond non-linearly to ambient weather, solar radiation and ventilation events [13, 14], and PV panels introduce further heterogeneity [15, 16, 17]. Hybrid, physics-based simulations capture those processes [18, 19, 20, 21, 22, 23] but are often too slow for on-line use; our results show that a graph-augmented neural network can bridge the gap when variable interactions are rich.
@@ 2 Data
This study draws on two high-frequency (15-min) data sets collected in Volos (Greece) under the REGACE project. GH2 spans an entire year (January to December 2020; 33 446 records) with six environmental variables, whereas GH4 is an autumn 2024 campaign (9 October to 20 November 2024; 5 000 records) that adds PAR and CO2 sensors.
GH2 (2020) – six-variable benchmark
@@
The inside-outside temperature trajectories for the autumn period in GH2 is shown in Figure 1. The winter and summer periods for GH2 are shown in 9 and 11 in Appendix A. During the summer period, the greenhouse employs a cooling system to prevent internal temperatures from rising excessively. Spring was excluded because of the high number of missing values. For modelling each season is split chronologically: the first 80 % of records form the training set, the remaining 20 % the test set.
Figure 1: GH2 temperatures inside and outside the greenhouse for the autumn period
@@
These physical relationships form the basis for the graph structure shown in Figure 2, where each edge represents a direct influence between variables. This directed graph approach allows our STGNN model to learn and weight these physical dependencies during the prediction process.
Figure 2: GH2 feature-interaction graph
GH4 (2024) – eight-variable greenhouse
Compared with GH2 the 2024 installation is instrumented with two additional physical dimensions: photosynthetically active radiation (PAR) and carbon-dioxide concentration, both measured inside and outside the structure:
@@
    (G4_PAR) / (OUT_PAR): PAR is the spectral band plants use for photosynthesis.
@@
    (G4_CO2) / (OUT_CO2): dry‐air CO2 mixing ratio (ppm).
Together with the common variables (air temperature, relative humidity and broadband solar radiation, these sensors yield five external drivers (OUT_Temp, OUT_Rad, OUT_PAR, OUT_CO2, OUT_RH) and three internal drivers (G4_PAR, G4_CO2, G4_Temp [target]). Data cover a single autumn window and are split chronologically 80/20 for modelling (train: 9 Oct 20:00–20 Nov 11:45; test: 20 Nov 12:00–30 Nov 21:45). Figure 3 shows the inside-outside temperature trajectories for the autumn period in GH4.
Figure 3: GH4 temperatures inside and outside the greenhouse for the autumn period
The rich PAR–CO2 couplings motivate the denser edge set of Figure 4, where directional links capture how radiation drives photosynthesis, which in turn modulates both internal CO2 and latent-heat–induced humidity changes.
Figure 4: GH4 feature-interaction graph
Missing-value treatment
GH2 contains 0–75 gaps per variable (winter: up to 68; autumn: 75). GH4 exhibits 273 missing points for every external variable and for G4_PAR / G4_CO2, and 274 for G4_Temp and G4_RH.
Gaps are imputed with a directional five-point mean: for a missing stamp at time $t$ (e.g. 13 Sep 01:00) we average the same quarter-hour on the two previous days ($t-24$ h, $t-48$ h) and the two following days ($t+24$ h, $t+48$ h). This approach ensured that the imputed values preserved the continuity in the time series without introducing noticeable distortions or flattening in the data sequences.
Normalisation and stationarity
All features are scaled to $[0,1]$ using min–max normalisation [24]. Augmented Dickey–Fuller tests confirm stationarity for every variable (all $p<0.01$), so no further differencing is applied [25, 26, 27].
@@ 3 Methodology
Neural Networks (NNs) are computational models inspired by the human brain’s neural structure. They consist of interconnected nodes (neurons) organized into layers: an input layer, one or more hidden layers, and an output layer. Each connection between neurons has an associated weight that is adjusted during training to minimize prediction error. Deep Neural Networks (DNNs) extend NNs by incorporating multiple hidden layers, enabling the modeling of complex, nonlinear relationships in data. Activation functions such as ReLU, sigmoid, and tanh introduce nonlinearity, while backpropagation adjusts the weights based on the error rate obtained in previous epochs [28].
@@ 3.1 Recurrent neural networks
Recurrent Neural Networks (RNNs) are specialized neural networks designed to process sequential data, making them highly suitable for time-series analysis [29]. Unlike traditional feedforward neural networks, RNNs have recurrent connections that allow information to persist across time steps, enabling the network to capture temporal dependencies.
@@
$\hat{y}_{t}=\bm{W}_{y}\bm{h}_{t}+\bm{b}_{y}$ (2)
where $\bm{W}_{h}$, $\bm{U}_{h}$, and $\bm{W}_{y}$ are weight matrices. $\bm{b}_{h}$ and $\bm{b}_{y}$ are bias vectors. $\tanh$ is the hyperbolic tangent activation function introducing nonlinearity. In deeper RNNs, multiple RNN layers are stacked, passing the hidden state from one layer to the next, allowing the model to capture more complex temporal patterns [30].
RNNs are effective for modeling how current and past environmental conditions influence future states, making them ideal for predicting variables like greenhouse temperature [31]. A known challenge with RNNs is the vanishing gradient problem, where gradients diminish during backpropagation through time, making it difficult to learn long-term dependencies [32]. However, since our data exhibits short-term dependencies, this issue do not impact our study [33].
@@ 3.1.1 RNN implementation
We begin by preparing the dataset, where each input sequence $\bm{X}^{(i)}$ consists of historical environmental measurements over a period of $T$ time steps. The target value $y^{(i)}$ is the greenhouse temperature at the next time step.
@@
The model is compiled using the Adam optimizer [34], which adapts the learning rate during training for efficient convergence. The loss function is set to Mean Squared Error (MSE), suitable for regression tasks.
@@ 3.2 Spatio-Temporal Graph Neural Networks
To model the spatio-temporal dynamics of the greenhouse microclimate, we employ directed STGNN. This approach integrates Graph Attention Networks (GATs) for spatial feature extraction [3] and Long Short-Term Memory (LSTM) networks [35] for temporal sequence modeling. Unlike traditional RNN-based models, which focus primarily on temporal sequences, STGNNs leverage Graph Neural Networks (GNNs) to incorporate topological and relational information among variables. This is well-aligned with recent advances in GNNs for time series [36]. By representing each environmental variable, such as internal temperature, humidity, and external factors like solar radiation and wind speed, as a node in a directed graph, we capture how changes in one variable may propagate to others through directional edges. Figure 2 and 4 illustrates this architecture. Each node corresponds to an environmental variable, while edges encode the direction of influence. At each time step $t$, node features $\bm{X}_{t}$ are extracted from the input sequence $\bm{X}^{(i)}$. To prevent data leakage, the feature of the target node (greenhouse temperature) at the current time step is masked (set to zero) before model input. The GAT layers learn to weigh the importance of edges and neighboring nodes, revealing spatial dependencies, while the LSTM layers capture long-term temporal patterns. More specifically, we adopt the GAT because its edge-wise attention coefficients are computed for each directed edge, allowing the model to weight asymmetric causal links (e.g., OUT_PAR $\rightarrow$ G4_PAR) that undirected GCN or GraphSAGE aggregators would blur. For the temporal branch we choose an LSTM cell: it offers the same gating mechanism as the Simple-RNN baseline (ensuring a fair comparison), remains numerically stable over our 96-step sequences, and is a well-established benchmark for reproducibility. It is important to emphasise that the graph is topological in feature space rather than in physical space: each node represents an environmental variable (e.g., PAR, CO2, humidity), not a sensor location, and edges encode functional or causal dependencies among variables rather than geographical proximity.